I need to find out if the two equations are perpendicular when placed on a graph. I cannot remember how to graph them. Thanks.
x+2y=5
2x+4y=9

Answer by mananth(16946)

x+2y=5...............1
2x+4y=9..............2
Eqyation1
when x= 0 y=2.5(0,2.5)
when y = 0 x= 5 ( 5 , 0 )
Equation 2
when x= 0 y= 2.25 ( 0 , 2.25 )
when y = 0 x= 4.5 ( 4.5 , 0 )

The lines have the same slope with different intercepts. So the lines are parallel